An object is moving so that its speed after t minutes is v (t ) = 1 + 4t + 3t2 meters per minute. How far does the object travel during the 3rd minute?
1
Expert's answer
2020-11-09T20:32:01-0500

path is the first derivative ofspeed. So we should find anintegral of the function of speed.


"\int_{2}^3 v(t)= 
=\int_{2}^3 (1+4t+3t^2)="

"=(t+4t^2\/2+3t^3\/3)=\n (t+2t^2+t^3)|_{2}^3 ="

"=3+18+27-2-8-8= 48-18= 30 m"

Answer: Diring 3rd minute object travel on 30 meters.